Find the product: (x + 1)(x^2 + x + 1)
Answer by stanbon(75887)

Find the product: (x + 1)(x^2 + x + 1)
----
= x(x^2+x+1)+1(x^2+x+1)
= x^3+x^2+x+x^2+x+1
= x^3+2x^2+2x +1
--------------
Cheers,
Stan H.
-------------